40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 9 17:34:11 2014, Nichin, Update, Electronics, PDFR: Beam pointing adjustments and DC measurements 
    Reply  Fri Jul 11 11:51:03 2014, Nichin, Update, Electronics, PDFR: List of DC transimpedances 
Message ID: 10166     Entry time: Wed Jul 9 17:34:11 2014     Reply to this: 10183
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: PDFR: Beam pointing adjustments and DC measurements 

 [Nichin, Manasa]

AIM: Taking DC output readings with multimeter for each PD to create a database (required for transimpedance calculations), by taking off the table tops. Also, making sure each PD is illuminated properly.

What we did:

  • In rack 1Y1: Diode laser controller was set to 150.0 mA at all times. This gave powers in the neighbourhood of 1mW at the end of fibers illuminating all PDs. The laser outputs light of 1064nm wavelength. The laser was switched off in the end.
  • Checked the collimation of the fiber for each PD. In some cases they were not focused to give a sharp spot, so we had to unmount the fibers and fix it and mount them back. Manasa did it initially and I learnt how it was done properly. Eventually I got better and did it myself (under her supervision)
  • Set the mount alignment for maximum illumination of the PD.
  • Record the power falling on the laser and also the DC voltage output. Any light that did not come from my fiber was blocked when taking the readings and then unblocked. I also took care of offset voltage present when taking the DC readings.

Recorded measurements:

REFL11:   Pinc = 0.91 mW         VDC = 34.9 mV 

REFL33:   Pinc = 0.83 mW         VDC = 33.2 mV 

REFL55:   Pinc = 1.08 mW         VDC = 42.7 mV 

REFL165: Pinc = 0.79 mW         VDC = 115.3 mV

AS55:         Pinc = 0.78 mW         VDC = 31.3 mV

POX11:      Pinc = 0.83 mW         VDC = 34.7 mV

POP22**:   Pinc = 1.08 mW         VDC = 5.82 mV

POY11:      Not illuminated; there was no optical fiber mount. Although, there was a fiber near it with a cap on the end. It also looks like there is no space to put in a new mount near the PD. 

REF PD:    Pinc = 1.19 mW         VDC = 8.2 V     (REF PD = New focus 1611)

**Note: The current POP 22 PD does not have 2 different outputs for DC and RF signals. I unplugged the RF cable from the output, took readings with the multimeter and then plugged back the RF cable.

Further work:

I will calculate the responsivity for each PD and compare it to the expected values. 

ELOG V3.1.3-